Weekly inspection – 11th May


Weather: big improvement with temperatures 17° C2 20° C. Set to continue with some showers

Forage: lots of chestnut, Hawthorn, cotton

Will need to see if blue colony needs intervention if the modified Snellgrove is not working

Also need to check on denim colony to see if the frame of eggs are being used for Queen rearing

14th May – Took Queen cell from denim and inserted into blue top box. For cyan took down to one Queen cell

Hive Record
CommentNext Inspection
AquaYY13 frames of bias – donated one to green (nuc) okay for spaceNothing in the super but is being workedAdded two frames to make up to full double broodCheck on swarming and super activity
BlueNNTop box looks to be queenless, bottom box, some some eggs and larvaePlenty in both brood boxesNeed to add Queen cell to top boxCheck on Queen present in bottom box, leave top box alone
CyanNN5 frames have capped brood, lots of QCs taken down to 21 kg in super, plenty of nectar coming inLooks like they rejected the queen, so will leave Queen cell in placeTake down to one Queen cell and use remaining elsewhere
DenimNNSmall amount of capped brewed and four sealed queen cells – taken down to 2Lots of nectar in brood box, so added superTook the donated frame of eggs from last week and started queen raisingCheck for emerged queen
FlaxYYFour frames of bias, large brood pattern but not many eggsPlenty of stores and lots of pollenReplace two monkey frames with foundationCheck if brood pattern expanding
GreenYYTwo frames of biasBringing in nectar so okayNeeded strengthening, so added capped brood frame from AquaCheck if continued strengthening required